OpenStudy (rcvelasco):

A 1 kg particle moves along an x axis, being propelled by a variable force directed along that axis. Its position is given by x = 1 m + (3 m/s)t + ct2 - (2 m/s3)t3 with x in meters and t in seconds.The factor c is a constant. At t = 3 s the force on the particle has a magnitude of 31 N and is in the negative direction of the axis. What is c?

OpenStudy (jiteshmeghwal9):

X=1+3t+ct^2-2t^3 Differentiating the expression with respect to time V=3+2ct-6t^2 Again differentiating with respect to time a=2c-12t

OpenStudy (jiteshmeghwal9):

Force=mass × acceleration F=1(2c-12t) At t=3 sec F=31N 31=2c-12(3)